OpenStudy (anonymous):

PLEASE HELP!!!!!! identify the vertex and the axis of symmetry for graph of the function y=3(x+2)^2-3 A. Vertex (2,-3) <---------- my answer not sure if its right axis of symmetry x=2 B vertex (-2,-3) axis of symmetry x=-2 C vertex (2,3) axis of symmetry x=2 D vertex (-2,3) axis of symmetry x=-2

OpenStudy (anonymous):

@ganeshie8 do you know this ?

ganeshie8 (ganeshie8):

vertec form :- \(\large y = a(x-h)^2 + k \) vertex = \( (h, k) \) axis of symmetry : \(x=h\)

ganeshie8 (ganeshie8):

\(\large y=3(x+2)^2-3 \) \(\large y=3(x- (-2))^2+ (-3) \)

ganeshie8 (ganeshie8):

compare it wid above, \(h = -2\) \(k = -3\) so... ?

OpenStudy (anonymous):

so it would end up being b and the axis of symmetry is -2?

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

thank you! :)

ganeshie8 (ganeshie8):

np :)